OpenStudy (anonymous):

You start out with $20 and then spend money in a store where every item is $3. Y = -3x + 20. What is the slope of this problem?

OpenStudy (anonymous):

Y = -3x + 20 =20-3x 20=money we have originally Y=money left 3=cost of item x= no. of item we buy therefore, slope = -3 which moeny spent on 1 item

OpenStudy (anonymous):

If that was the slop can you tell me the y intercept?

OpenStudy (anonymous):

put x=0, we can find y-intercept which is 20 20 is the money we have originally

OpenStudy (anonymous):

Memorize the linear formula: y = mx + b ( m is slope, b is y intercept) Match them: Y = -3x + 20

OpenStudy (anonymous):

@gageallen15 Of course CoCoTsoi's explanation is far better than mine, since it implies real concept. Just try to show you the shortcut in case :P

OpenStudy (anonymous):

Thank you :)